# Doubly-Dynamic ISAC Precoding for Vehicular Networks: A Constrained Deep Reinforcement Learning (CDRL) Approach

### v1

@@ Abstract
Integrated sensing and communication (ISAC) technology is essential for enabling the vehicular networks. However, the communication channel in this scenario exhibits time-varying characteristics, and the potential targets may move rapidly, creating a doubly-dynamic phenomenon. This nature poses a challenge for real-time precoder design. While optimization-based solutions are widely researched, they are complex and heavily rely on perfect prior information, which is impractical in double dynamics. To address this challenge, we propose using constrained deep reinforcement learning (CDRL) to facilitate dynamic updates to the ISAC precoder design. Additionally, the primal dual-deep deterministic policy gradient (PD-DDPG) and Wolpertinger architecture are tailored to efficiently train the algorithm under complex constraints and variable numbers of users. The proposed scheme not only adapts to the dynamics based on observations but also leverages environmental information to enhance performance and reduce complexity. Its superiority over existing candidates has been validated through experiments.
@@ I Introduction
The Internet of Vehicles (IoV) is essential for supporting an efficient and reliable intelligent transportation system, with high-speed communications and high-precision sensing serving as cornerstones [1, 2]. In comparison to a separate communication and sensing system, integrated sensing and communication (ISAC) offers the potential for mutual enhancement, making it a recent research focus in vehicular networks [3, 4, 5]. However, the dynamic nature of the road environment introduces time-varying characteristics in the wireless channel due to user and scatterer movements. Simultaneously, the targets themselves move rapidly and unpredictably, presenting a doubly-dynamic challenge for ISAC.
@@
In an effort to mitigate the complexity issue and lower the dependence on real-time prior information, methods such as message passing (MP) and particle filter (PF) with modified operating protocols have been proposed in [10, 11]. However, these solutions still struggle to address the doubly-dynamics challenge, since the mutual interference between communication and sensing can degrade the performance of these algorithms. Furthermore, as the number of users increases, the complexity of these methods increase significantly, diminishing the timeliness of ISAC systems.
Reinforcement Learning (RL), known for its adaptability to dynamics without prior assumptions, has garnered attention in applications to ISAC systems in [12, 13]. However, applying RL to the considered doubly-dynamic problem still faces two main difficulties: Firstly, existing RL algorithms struggle to handle complex constraints in this problem. Secondly, the parameter tuning in precoding design entails a large action space, the size of which may vary with the number of users, posing difficulty in action selection and policy updates.
@@ II-A Transmitted Waveform
where $P_{\text{t}}$ denotes the transmitting power, $\bm{s}_{\text{m}}^{(t,l)}=\left[s_{\text{m},1}^{(t,l)},\cdots,s_{\text{m},\text{U}}^{(t,l)}\right]^{\text{T}}$ is the communication symbols for the $U$ users at the $m$-th subcarrier with $\mathbb{E}\left[\bm{s}_{\text{m}}\bm{s}_{\text{m}}^{\text{H}}\right]=\bm{I}_{\text{U}}$, and $\bm{F}^{(t)}\in\mathbb{R}^{N_{\text{t}}\times\text{U}}$ is the precoder at each subcarrier, updated at a subframe-level timescale. Since the analog precoder is composed of phase shifters between radio frequency (RF) chains and the antenna array, each element of $\bm{F}$ has constant modulus, i.e., $|[\bm{F}]_{i,j}|^{2}=1$.
@@ II-B Time-varying Communication Model
Figure 1: An illustration of ISAC system in doubly-dynamic scenarios.
@@ II-C Time-varying Sensing Model
Denote $\left[\theta_{\text{s}}^{t},d_{\text{s}}^{t}\right]$ is the target’s position in polar domain in the $t$-th subframe. Define $\bm{G}^{(t)}_{\text{m}}=\alpha^{(t)}_{\text{m}}\bm{a}_{r}(\theta_{\text{s}}^{t})\bm{a}_{\text{t}}^{\text{H}}(\theta_{\text{s}}^{t})$ to be the target response matrix at the $m$-th subcarrier, with $\alpha_{\text{m}}^{(t)}$ being the complex reflection coefficient that incorporates the round-trip path loss, the radar cross-section (RCS) and the Doppler of the target. Accordingly, the received echo becomes
@@ III Optimization-based ISAC precoding
From communication-centric perspective, the objective of the doubly-dynamic ISAC precoding is to minimize the averaged CRLB for angle estimation, while also ensuring SE throughout the frames, and taking into account the constant-modulus constraint of the precoder. Therefore, the corresponding problem can be formulated as
@@
$~{}|[\bm{F}^{(t)}]_{i,j}|^{2}=1.\vspace{-0.5cm}$ (9b)
Solving this optimization problem directly can be tricky due to the complex objective function and constant-modulus constraints. Furthermore, the ICI induced by Doppler imposes a more complex constraint in (9a) than in [4]. To address these challenges, we approach the sequential optimization on a subframe-by-subframe basis as
@@
Subsequently we employ the semi-definite relaxation (SDR) technique to iteratively derive a sub-optimal solution at $t$. Denote $\bm{W}_{\text{u}}=\bm{f}_{\text{u}}\bm{f}_{\text{u}}^{\text{H}}$ and $\bm{Q}_{\text{m},\text{k},\text{u}}=\bm{h}_{\text{m},\text{k},\text{u}}^{\text{H}}\bm{h}_{\text{m},\text{k},\text{u}}$. Then we reformulate the objective and constraints as
@@
Discussion on the applicability: This optimization-based method solves the problem at the cost of high computational complexity, scaling up to $\mathcal{O}(N_{\text{t}}^{3.5}\log(1/\epsilon))$, given a solution accuracy $\epsilon$. This results in an unacceptable delay, especially when the antenna array becomes large, as is the case in the studied scenario. Furthermore, in the context of the double dynamics, obtaining prior information $\bm{H}^{(t)}$ and $\theta_{\text{s}}^{(t)}$ for optimizing $\bm{F}^{(t)}$ is not readily available. Consequently, the optimization-based method may be no longer applicable.
@@ IV-A Constrained Markov Decision Process Formulation
The doubly-dynamic scenario is initially modeled as a CMDP. Once the environment state is observed in subframe $t$, BS takes an action according to its strategy. Subsequently, BS receives a reward as well as a cost, and the environment evolves to the next state with a certain transition probability. This process continues, with the new state being observed at each subframe, and new actions are taken accordingly. The essential elements of this CMDP are described below in detail.
@@ IV-A1 State Space
$\bm{P}^{(t)}(n_{\text{x}},\!n_{\text{y}})\!=\!\!\begin{cases}1,\!\!&\text{if $\{[\theta_{n_{\text{x}}},\!\theta_{n_{\text{x}}\!+\!1})\!\times\![d_{n_{\text{y}}},\!d_{n_{\text{y}}\!+\!1})\}\!\cap\mathcal{U}^{(t)}\!\!\neq\!\emptyset$}\\ 2,\!\!&\text{if $\{[\theta_{n_{\text{x}}},\!\theta_{n_{\text{x}}\!+\!1})\!\times\![d_{n_{\text{y}}},\!d_{n_{\text{y}}\!+\!1})\}\!\cap\mathcal{T}^{(t)}\!\!\neq\!\emptyset$}\\ 0,\!\!&\text{Otherwise},\end{cases}\vspace{-0.2cm}$
@@ IV-A3 Reward and cost functions
$r^{(t)}=-\text{CRLB}(\theta_{\text{s}}^{(t)}),~{}~{}c^{(t)}=-\text{SE}^{(t)}.\vspace{-0.2cm}$ (12)
@@
$R(\mu)\!=\!\mathbb{E}_{\mu}\!\left[\!\sum_{t=t_{0}}^{T}\gamma^{t-t_{0}}r^{(t)}\!\right],~{}C(\mu)\!=\!\mathbb{E}_{\mu}\!\left[\!\sum_{t=t_{0}}^{T}\gamma^{t-t_{0}}c^{(t)}\!\right],$ (13)
@@ IV-A4 State transition
In the $(t\!+\!1)$-th subframe, the state $\bm{s}^{(t+1)}\!\!=\!\!\{\bm{S}_{\text{H}}^{(t+1)},\!\bm{S}_{\text{P}}^{(t+1)}\}$ is evolved as $\bm{S}_{\text{H}}^{(t+1)}=\bm{S}_{\text{H}}^{(t)}$, and $\bm{S}_{\text{P}}^{(t+1)}=\left[\bm{P}^{(t-2)},\bm{P}^{(t-1)},\bm{P}^{(t)}\right]$. The channel-related state remains unaltered, while the position-related state can be refreshed with the latest estimation, $\bm{P}^{(t)}$, through the BS-end echo processing at $t$. The evolution of the position spectrum $\bm{P}^{(t)}$ across subframes is contingent upon the real positions of the users/scatterers, and the target, which are parameterized by [$\theta^{(t)}_{\text{u},\text{p}}$, $\tau^{(t)}_{\text{u},\text{p}}$] and [$\theta^{(t)}_{\text{s}}$, $d^{(t)}_{\text{s}}$] respectively. These parameters vary according to respective velocities $\bm{v}_{\text{u},\text{p}}$ and $\bm{v}_{\text{s}}$, where $|\bm{v}_{\text{u},\text{p}}|,|\bm{v}_{\text{s}}|\!\sim\!\mathcal{U}[v_{\text{min}},v_{\text{max}}]$ with the angle randomly selected from 0 and $\pi$. The specific evolution process characterized by kinematic equations can refer to [10].
@@ IV-B1 Learning architecture
The reward and cost critic networks, parameterized by $\phi_{\text{R}}$ and $\phi_{\text{C}}$, map from the joint state space and action space to reward and cost respectively, playing a role of action evaluation: $Q_{\text{R}}(\bm{s}^{(t)},\bm{a}^{(t)}|\phi_{\text{R}})$ and $Q_{\text{C}}(\bm{s}^{(t)},\bm{a}^{(t)}|\phi_{\text{C}})$. The actor network $\mu$, parameterized by $\phi_{\text{A}}$, maps from the state space to an action $\bm{a}\in\mathcal{R}^{N_{\text{A}}}$ as $\mu(\bm{s}^{(t)}|\phi_{\text{A}})=\bm{a}^{(t)}$. To reduce fluctuations in target values during training and expedite convergence, we employ target networks $Q_{\text{R}}^{\prime}$, $Q_{\text{C}}^{\prime}$ and $\mu^{\prime}$ parameterized by $\phi_{\text{R}}^{\prime}$, $\phi_{\text{C}}^{\prime}$ and $\phi_{\text{A}}^{\prime}$ respectively. To solve the considered CMDP in (15), we first adopt the Lagrangian relaxation procedure to transform the problem in (14) to an unconstrained one as
@@ IV-B3 Primal-Dual Network Update
To solve the unconstrained minimax problem in (15), we alternately update the policy parameterized by $\phi_{\text{A}}$ and the dual variable $\lambda$ in each iteration, while $\phi_{\text{R}}$ and $\phi_{\text{C}}$ are updated for more precise policy evaluation. To be specific, a batch of $N_{\text{b}}$ transition tuples are randomly sampled from $\mathcal{M}$, and the target reward and target cost is calculated as
@@
3:Initialize replay buffer $\mathcal{M}=\emptyset$, dual variation $\lambda=0$, mini-batch size $N_{\text{b}}=64$ and discount factor $\gamma=0.6$;
@@ V Experiments
In this section, simulation results are presented to evaluate the performance of the proposed scheme. Unless otherwise specified, the system parameters are set as: $N_{\text{t}}=N_{\text{r}}=32$, $M=32$, $L_{\text{cp}}=8$, $L=32$, $T=100$, $U^{\text{max}}=16$, $N_{\text{B}}=32$, $G_{\text{t}}=64$. The central frequency point is set as $f_{\text{c}}=28$ GHz with interval $\Delta f=30$ kHz. $T_{\text{s}}=\frac{1}{\Delta f}=33.3~{}\mu$s. $g_{\text{rp}}(\cdot)$ is set as the raised-cosine filter with roll-off factor as 0.4. The channel-related parameters are set as $N_{\text{d}}=8$, $P_{\text{u}}=4$, $\sigma_{\beta}=1$. $\alpha_{\text{m}}$ has a random phase uniformly distributed in $[0,2\pi)$, and its amplitude obeys that $|\alpha_{\text{m}}|(\text{dB})=5\log_{10}(\sigma_{\text{RCS}})-10\log_{10}(f_{\text{m}})-20\log_{10}(d)+110$ according to [19], where $\sigma_{\text{RCS}}=10$ is the RCS of the target. The noise variances are set as $\sigma_{\text{c}}^{2}=\sigma_{\text{z}}^{2}=-10$ dBm. SNR is defined as $\frac{P_{\text{t}}}{\sigma_{\text{c}}^{2}}$. CDRL-PC and CDRL-PO denote the proposed algorithm with different state input respectively (positions plus initial CSI, positions-only). The perfect prior means that the instantaneous CSI and precise positions are available at BS. Main benchmarks include the optimization-based method in Sec. III, greedy selection in [20] and random selection. In the experiment environment, the sensing range is $[-\pi,\pi)\times[0~{}\text{m},50~{}\text{m})$. The velocity (m/s) of the target, the users and the related scatters obey a uniform distribution of $\mathcal{U}[10,30]$. Main hyper-parameters for model training in experiments are set as: the size of the replay buffer $\mathcal{M}$ is 1024, the batch size is $N_{\text{b}}=64$, the learning rates of the actor, reward critic and cost critic are $(\kappa_{\text{A}},\kappa_{\text{R}},\kappa_{\text{C}})=(0.05,0.05,0.1)$. The discounting factor is $\gamma=0.6$, the parameters of the O-U noise for exploration is $(\mu_{\text{OU}},\theta_{\text{OU}},\sigma_{\text{OU}})=(0,0.5,0.3)$. We adopt Adam as the optimizer. The proposed algorithm is developed using PyTorch and on a Window™-based machine running an NVIDIA™ GeForce RTX4060 GPU.
@@
Figure 3: Cumulative reward and cost versus epoches.

### v2

@@ Abstract
Integrated sensing and communication (ISAC) technology is essential for enabling vehicular networks. However, the communication channel in this scenario exhibits time-varying characteristics, and the potential targets may move rapidly, creating a doubly-dynamic phenomenon. This nature poses a challenge for real-time precoder design. While optimization-based solutions are widely researched, they are complex and heavily rely on perfect channel-related information, which is impractical in double dynamics. To address this challenge, we propose using constrained deep reinforcement learning (CDRL) to facilitate dynamic updates to the ISAC precoder. Additionally, the primal dual-deep deterministic policy gradient (PD-DDPG) and Wolpertinger architecture are tailored to efficiently train the algorithm under complex constraints and varying numbers of users. The proposed scheme not only adapts to the dynamics based on observations but also leverages environmental information to enhance performance and reduce complexity. Its superiority over existing candidates has been validated through experiments.
@@ I Introduction
The Internet of Vehicles (IoV) is essential for supporting an efficient and reliable intelligent transportation system, with high-speed communications and high-precision sensing serving as cornerstones [1, 2]. In comparison to separate communication and sensing systems, integrated sensing and communication (ISAC) offers the potential for mutual enhancement, making it a recent research focus in vehicular networks [3, 4, 5, 6]. However, the dynamic nature of the road environment introduces time-varying characteristics in the wireless channel due to user and scatterer movements. Simultaneously, the targets themselves move rapidly and unpredictably, presenting a doubly-dynamic challenge for ISAC.
@@
In an effort to mitigate the complexity issue and lower the dependence on real-time prior information, methods such as message passing (MP) and particle filter (PF) with modified operating protocols have been proposed in [11, 12]. However, these solutions still struggle to address the doubly-dynamics challenge, since the mutual interference between communication and sensing can degrade the performance of these algorithms. Furthermore, as the number of users increases, the complexity of these methods increases significantly, diminishing the timeliness of ISAC systems.
Reinforcement Learning (RL), known for its adaptability to dynamics without prior assumptions, has garnered attention in applications to ISAC systems [13, 14]. However, applying RL to the considered doubly-dynamic problem still faces two main difficulties: Firstly, existing RL algorithms in [15, 16] struggle to handle complex constraints in this problem, resulting in performance degradation. Secondly, the parameter tuning in precoding design entails a large action space, the size of which may vary with the number of users, posing difficulty in action selection and policy updates.
@@ II-A Transmitted Waveform
where $P_{\text{t}}$ denotes the transmitting power, $\bm{s}_{\text{m}}^{(t,l)}=\left[s_{\text{m},1}^{(t,l)},\cdots,s_{\text{m},\text{U}}^{(t,l)}\right]^{\text{T}}$ contains the communication symbols for $U$ users at the $m$-th subcarrier with $\mathbb{E}\left[\bm{s}_{\text{m}}\bm{s}_{\text{m}}^{\text{H}}\right]=\bm{I}_{\text{U}}$, and $\bm{F}^{(t)}\in\mathbb{R}^{N_{\text{t}}\times\text{U}}$ is the precoder at each subcarrier, updated at a subframe-level timescale. Since the analog precoder is composed of phase shifters between radio frequency (RF) chains and the antenna array, each element of $\bm{F}$ has constant modulus, i.e., $|[\bm{F}]_{i,j}|^{2}=1$.
@@ II-B Time-varying Communication Model
Figure 1: An illustration of the ISAC system in doubly-dynamic scenarios.
@@ II-C Time-varying Sensing Model
Denote $\left[\theta_{\text{s}}^{t},d_{\text{s}}^{t}\right]$ is the target’s position in the polar domain in the $t$-th subframe. Define $\bm{G}^{(t)}_{\text{m}}=\alpha^{(t)}_{\text{m}}\bm{a}_{r}(\theta_{\text{s}}^{t})\bm{a}_{\text{t}}^{\text{H}}(\theta_{\text{s}}^{t})$ to be the target response matrix at the $m$-th subcarrier, with $\alpha_{\text{m}}^{(t)}$ being the complex reflection coefficient that incorporates the round-trip path loss, the radar cross-section (RCS) and the Doppler of the target. Accordingly, the received echo becomes
@@ III Optimization-based ISAC precoding
From the communication-centric perspective, the objective of the doubly-dynamic ISAC precoding is to minimize the averaged CRLB for angle estimation, while also ensuring SE throughout the frames, and taking into account the constant-modulus constraint of the precoder. Therefore, the corresponding problem can be formulated as
@@
$~{}|[\bm{F}^{(t)}]_{i,j}|^{2}=1,\vspace{-0.5cm}$ (9b)
given the perfect $\overline{\bm{H}}$ and perfect $\bm{G}_{\text{m}}$ in each subframe. Solving this optimization problem directly can be tricky due to the complex objective function and constant-modulus constraints. Furthermore, the ICI induced by Doppler imposes a more complex constraint in (9a) than in [4]. To address these challenges, we approach the sequential optimization on a subframe-by-subframe basis as
@@
Subsequently, we employ the semi-definite relaxation (SDR) technique to iteratively derive a sub-optimal solution at $t$. Denote $\bm{W}_{\text{u}}=\bm{f}_{\text{u}}\bm{f}_{\text{u}}^{\text{H}}$ and $\bm{Q}_{\text{m},\text{k},\text{u}}=\bm{h}_{\text{m},\text{k},\text{u}}^{\text{H}}\bm{h}_{\text{m},\text{k},\text{u}}$. Then we reformulate the objective and constraints as
@@
Discussion on the applicability: This optimization-based method solves the problem at the cost of high computational complexity, scaling up to $\mathcal{O}(N_{\text{t}}^{3.5}\log(1/\epsilon))$, given a solution accuracy $\epsilon$. This results in an unacceptable delay, especially when the antenna array becomes large, as is the case in the studied scenario. Furthermore, in the context of the double dynamics, obtaining prior information $\bm{H}^{(t)}$ and $\theta_{\text{s}}^{(t)}$ for optimizing $\bm{F}^{(t)}$ is not readily available. Consequently, the optimization-based method may no longer be applicable.
@@ IV-A Constrained Markov Decision Process Formulation
Due to the evolving characteristics of wireless channels and positions, the doubly-dynamic scenario is initially modeled as a CMDP. Once the environment state is observed in subframe $t$, BS takes an action according to its strategy. Subsequently, BS receives a reward as well as a cost, and the environment evolves to the next state with a certain transition probability. This process continues, with the new state being observed at each subframe, and new actions are taken accordingly. The essential elements of this CMDP are described below in detail.
@@ IV-A1 State Space
$\bm{P}^{(t)}(n_{\text{x}},\!n_{\text{y}})\!=\!\!\begin{cases}1,\!\!&\text{if $\{[\theta_{n_{\text{x}}},\!\theta_{n_{\text{x}}\!+\!1})\!\times\![d_{n_{\text{y}}},\!d_{n_{\text{y}}\!+\!1})\}\!\cap\mathcal{U}^{(t)}\!\!\neq\!\varnothing$},\\ 2,\!\!&\text{if $\{[\theta_{n_{\text{x}}},\!\theta_{n_{\text{x}}\!+\!1})\!\times\![d_{n_{\text{y}}},\!d_{n_{\text{y}}\!+\!1})\}\!\cap\mathcal{T}^{(t)}\!\!\neq\!\varnothing$},\\ 0,\!\!&\text{Otherwise},\end{cases}\vspace{-0.2cm}$
@@ IV-A3 Reward and cost functions
$r^{(t)}=-\text{CRLB}(\theta_{\text{s}}^{(t)}),~{}~{}c^{(t)}=-\text{SE}^{(t)}.\vspace{-0.1cm}$ (12)
@@
$R(\mu)\!=\!\mathbb{E}_{\mu}\!\left[\!\sum_{t=t_{0}}^{T}\gamma^{t\!-\!t_{0}}r^{(t)}\!\right],~{}C(\mu)\!=\!\mathbb{E}_{\mu}\!\left[\!\sum_{t=t_{0}}^{T}\gamma^{t\!-\!t_{0}}c^{(t)}\!\right],\vspace{-0.15cm}$ (13)
@@ IV-A4 State transition
In the $(t\!+\!1)$-th subframe, the state $\bm{s}^{(t+1)}\!\!=\!\!\{\bm{S}_{\text{H}}^{(t+1)},\!\bm{S}_{\text{P}}^{(t+1)}\}$ is evolved as $\bm{S}_{\text{H}}^{(t+1)}=\bm{S}_{\text{H}}^{(t)}$, and $\bm{S}_{\text{P}}^{(t+1)}=\left[\bm{P}^{(t-2)},\bm{P}^{(t-1)},\bm{P}^{(t)}\right]$. The channel-related state remains unaltered, while the position-related state can be refreshed with the latest estimation, $\bm{P}^{(t)}$, through the BS-end echo processing at $t$. The evolution of the position spectrum $\bm{P}^{(t)}$ across subframes is contingent upon the real positions of the users/scatterers, and the target, which are parameterized by [$\theta^{(t)}_{\text{u},\text{p}}$, $\tau^{(t)}_{\text{u},\text{p}}$] and [$\theta^{(t)}_{\text{s}}$, $d^{(t)}_{\text{s}}$] respectively. These parameters vary according to respective velocities $\bm{v}_{\text{u},\text{p}}$ and $\bm{v}_{\text{s}}$, where $|\bm{v}_{\text{u},\text{p}}|,|\bm{v}_{\text{s}}|\!\sim\!\mathcal{U}[v_{\text{min}},v_{\text{max}}]$ with the angle randomly selected from 0 and $\pi$. The specific evolution process characterized by kinematic equations can be referred to [11].
@@ IV-B1 Learning architecture
To tackle this constrained optimization problem, the reward and cost critic networks, parameterized by $\phi_{\text{R}}$ and $\phi_{\text{C}}$, map from the joint state space and action space to reward and cost respectively, playing a role of action evaluation: $Q_{\text{R}}(\bm{s}^{(t)},\bm{a}^{(t)}|\phi_{\text{R}})$ and $Q_{\text{C}}(\bm{s}^{(t)},\bm{a}^{(t)}|\phi_{\text{C}})$. The actor network $\mu$, parameterized by $\phi_{\text{A}}$, maps from the state space to an action $\bm{a}\in\mathbb{R}^{N_{\text{A}}}$ as $\mu(\bm{s}^{(t)}|\phi_{\text{A}})=\bm{a}^{(t)}$. To reduce fluctuations in target values during training and expedite convergence, we employ target networks $Q_{\text{R}}^{\prime}$, $Q_{\text{C}}^{\prime}$ and $\mu^{\prime}$ parameterized by $\phi_{\text{R}}^{\prime}$, $\phi_{\text{C}}^{\prime}$ and $\phi_{\text{A}}^{\prime}$ respectively. To solve the considered CMDP in (15), we first adopt the Lagrangian relaxation procedure to transform the problem in (14) to an unconstrained one as
@@ IV-B3 Primal-Dual Network Update
To facilitate the model training, we adopt a primal dual-deep deterministic policy gradient (PD-DDPG) procedure, where the policy parameterized by $\phi_{\text{A}}$ and the dual variable $\lambda$ are updated alternately for objective maximization while the critics $\phi_{\text{R}}$ and $\phi_{\text{C}}$ are updated for more precise policy evaluation. To be specific, a batch of $N_{\text{b}}$ transition tuples are randomly sampled from $\mathcal{M}$, and the target reward and target cost is calculated as
@@
3:Initialize replay buffer $\mathcal{M}=\varnothing$, dual variation $\lambda=0$, mini-batch size $N_{\text{b}}=64$ and discount factor $\gamma=0.6$;
@@ V Experiments
In this section, simulation results are presented to evaluate the performance of the proposed scheme. The adopted dataset incorporates both time-varying channel data and moving target information to reflect practical double dynamics [21]. Unless otherwise specified, the system parameters are set as: $N_{\text{t}}=N_{\text{r}}=32$, $M=32$, $L_{\text{cp}}=8$, $L=32$, $T=100$, $U^{\text{max}}=16$, $N_{\text{B}}=32$, $G_{\text{t}}=64$. The central frequency point is set as $f_{\text{c}}=28$ GHz with interval $\Delta f=30$ kHz. $T_{\text{s}}=\frac{1}{\Delta f}=33.3~{}\mu$s. $g_{\text{rp}}(\cdot)$ is set as the raised-cosine filter with roll-off factor as 0.4. The channel-related parameters are set as $N_{\text{d}}=8$, $P_{\text{u}}=4$, $\sigma_{\beta}=1$. $\alpha_{\text{m}}$ has a random phase uniformly distributed in $[0,2\pi)$, and its amplitude obeys that $|\alpha_{\text{m}}|(\text{dB})=5\log_{10}(\sigma_{\text{RCS}})-10\log_{10}(f_{\text{m}})-20\log_{10}(d)+110$ according to [22], where $\sigma_{\text{RCS}}=10$ is the RCS of the target. The noise variances are set as $\sigma_{\text{c}}^{2}=\sigma_{\text{z}}^{2}=-10$ dBm. SNR is defined as $\frac{P_{\text{t}}}{\sigma_{\text{c}}^{2}}$. In the experiment environment, the sensing range is $[-\pi,\pi)\times[0~{}\text{m},50~{}\text{m})$. The velocity (m/s) of the target, the users and the related scatters obey a uniform distribution of $\mathcal{U}[10,30]$. Main hyper-parameters for model training in experiments are set as: the size of the replay buffer $\mathcal{M}$ is $10^{5}$, the batch size is $N_{\text{b}}=64$, the learning rates of the actor, reward critic and cost critic are $(\kappa_{\text{A}},\kappa_{\text{R}},\kappa_{\text{C}})=(0.05,0.05,0.1)$. The discounting factor is $\gamma=0.6$, the parameters of the O-U noise for exploration is $(\mu_{\text{OU}},\theta_{\text{OU}},\sigma_{\text{OU}})=(0,0.5,0.3)$. We adopt Adam as the optimizer. CDRL-PC and CDRL-PO denote the proposed algorithm with different state inputs respectively (positions plus initial CSI, positions-only). The perfect prior means that the instantaneous CSI and precise positions are available at BS. Main benchmarks include the optimization-based method in Sec. III, greedy selection in [23] and random selection. The proposed algorithm is developed using PyTorch and on a Window™-based machine running an NVIDIA™ GeForce RTX4060 GPU.
@@
Figure 3: Cumulative reward and cost versus epochs.